OpenStudy (anonymous):

PLEASE HELP!!!! find the exact value of arctan(tan[5π/4])

OpenStudy (jdoe0001):

an inverse trig function, will give you the ANGLE that has that value say what... \(\bf cos^{-1}(0)\quad \textit{what angle has a cosine of 0?, well }90^o\ or \ 270^o\) so that means that if I'd do say... \(\bf cos(90^o)=0\ or \ cos(270^o) = 0\\ \quad \\ so\quad cos^{-1}[cos(90^o)]=90^o\qquad thus\\ \quad \\ cos^{-1}[cos(\theta)]=\theta\qquad sin^{-1}[sin(\theta)]=\theta\qquad tan^{-1}[tan(\theta)]=\theta\)

OpenStudy (jdoe0001):

so what do you think is the ANGLE that will be spit out by the inverse function \(\bf tan^{-1}\left[tan\left(\frac{5\pi}{4}\right)\right]\quad ?\)

OpenStudy (anonymous):

lol